OpenStudy (anonymous):

How do I make (k/2)(5k-1) + 5(k+1)-3 equal to ((k+1)/2)(5k+4)?

OpenStudy (anonymous):

by multiplying both part of the equation by 2

OpenStudy (anonymous):

[k/2(5k-1)+5(k+1)-3 = (k+1)/2 (5k+4)\]

OpenStudy (anonymous):

How do I simplify the left side?

OpenStudy (anonymous):

before and after you simplify it, it doesn't matter. . to solve, use only distributive property of multiplication. . . .

OpenStudy (anonymous):

I need to make the left side look like the right side

jhonyy9 (jhonyy9):

k(5k-1)+10(k+1)-6=(k+1)(5k+4) 5k2-k+10k+10-6=5k2+4k+5k+4 5k2+9k+4=5k2+9k+4 0=0

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

do yo understand it ?

OpenStudy (anonymous):

yeah you got rid of 2 by multipluying it, right >?

jhonyy9 (jhonyy9):

you need multiply by 2 because 2 is denominator common

OpenStudy (anonymous):

ok, thats the part i had trouble with

OpenStudy (anonymous):

that's what i'm talking about thanks jhonny for expounding it. . .

jhonyy9 (jhonyy9):

after multiply term by term both sides and get that the left side is equal right side

jhonyy9 (jhonyy9):

yw jerwyn

OpenStudy (anonymous):

thanks again

jhonyy9 (jhonyy9):

yw

jhonyy9 (jhonyy9):

medal ?

OpenStudy (anonymous):

done

jhonyy9 (jhonyy9):

ok good luck bye

OpenStudy (anonymous):

bye